OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. A 24-foot beam is to be cut in two. If one portion is half the length of the other portion, what are the two lengths of the portions?

OpenStudy (anonymous):

it'll be 8 and 16

OpenStudy (anonymous):

how did you get that if 12 is half of 24

OpenStudy (anonymous):

you get it by 24=.5x+x and solve for x and you get 16. so .5(16)=8 and the other part is x=16

OpenStudy (anonymous):

if you add 8 and 16 together you get the full length of 24.

OpenStudy (gw2011):

Let x=1 portion Let (1/2)x=the other portion x+(1/2)x=24 (3/2)x=24 x=(24)(2/3) x=16 which is portion 1 (1/2)x=8 which is the other portion

OpenStudy (anonymous):

so I was right :D..Lol